OpenStudy (anonymous):

What is the value of 2^log2(7)

OpenStudy (anonymous):

7

OpenStudy (zarkon):

\[a^{\log_{a}(b)}=b\]

OpenStudy (anonymous):

how do you do a subscript in the logarithm on a calculator?

OpenStudy (zarkon):

\[log_{a}(b)=\frac{\ln(b)}{\ln(a)}\]

OpenStudy (anonymous):

oh oh oh ok thank you